OpenStudy (anonymous):

x + 2y = 5 3x + 5y = 14 Solve the system of equations. (3, 1) (7, -1) (2, 3/2)

OpenStudy (anonymous):

the easiest way to do this problem is like this.... x+2y=5 x=-2y+5 then substituite this in for x in the other equation 3(-2y+5) +5y = 14 then distribute... -6y+15+5y=14 combine like terms... -y+15=14 -y=-1 divide by negative one gettin y = 1 substitue 1 into this equation... x=2(1)+5 x=7 y=1